OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: Chart with two rows and four columns. The first row is labeled Number of Pages and contains the number fifty in the first column, one hundred in the second column, one hundred fifty in the third column, and two hundred in the fourth column. The second row is labeled Cost and contains the numbers one point two five in the first column, one point eight five in the second column, two point four five in the third column, and three point zero five in the fourth column

OpenStudy (anonymous):

A. y = 0.012x + 49.985 B. y = 83.3x + 0.65 C. y = 0.012x + 1.25 D. y = 0.012x + 0.65

OpenStudy (k12girlohva):

http://brainly.com/question/254032

OpenStudy (anonymous):

is this your chart:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You know, the fastest and easiest way to solve this is to simply plug these values in to a Microsoft Excel spreadsheet, and then have it calculate the linear regression of the plotted points. Of course, you may calculate it manually yourself. If that's the case, I'd recommend getting familiar with the linear regression formula first. Good luck!